### v1

# Barycenter curvature-dimension condition for extended metric measure spaces

Abstract
In this survey, we introduce a new curvature-dimension condition for extended metric measure spaces, called Barycenter-Curvature Dimension condition BCD, from the perspective of Wasserstein barycenter.
Keywords: Wasserstein barycenter, metric measure space, curvature-dimension condition, Ricci curvature, optimal transport
MSC 2020: 53C23, 51F99, 49Q22
1 Optimal transport and displacement convexity
The theory of optimal transport stems from a practical problem:
What is the most economical way to transport materials from one place to another, with given supply and demand?
In the language of mathematics, this problem is to minimize the following quantity:
$I[\pi]=\int_{X\times X}c(x,y)\,{\mathrm{d}}\pi(x,y)~{}~{}~{}\text{for }\pi\in\Pi(\mu,\nu),$
where $c:X\times X\rightarrow[0,\infty]$ is a cost function and $\Pi(\mu,\nu)$ is composed of all probability measure (transport plan) $\pi$ on $X\times X$ whose marginals are $\mu$ and $\nu$. We denote the infimum of $I[\pi]$ by $T_{c}(\mu,\nu)$.
Let $(X,\mathsf{d})$ be a metric space and consider the optimal transport problem with the cost function $c=\mathsf{d}^{2}$. We define the Wasserstein distance $W_{2}$ as $W_{2}^{2}(\mu,\nu):=T_{\mathsf{d}^{2}}(\mu,\nu)$. Evidently, it is an extended metric (allows distance $W_{2}=\infty$) on the space of Borel probability measure $\mathcal{P}(X)$. To make it become a metric, we often restrict the study on subspace $\mathcal{P}_{2}(X,\mathsf{d}):=\left\{\mu\in\mathcal{P}(X):\int_{X}\mathsf{d}^{2}(x,x_{0})\,{\mathrm{d}}x<\infty~{}\text{for some}~{}x_{0}\in X\right\}$. The metric space $(\mathcal{P}_{2}(X,\mathsf{d}),W_{2})$ is called Wasserstein space. The geometry theory of Wasserstein space has a wide range of applications in many fields, including differential equations, geometric and functional inequalities. We refer to [Vil09] for more materials. In this survey, we mainly focus on the connection between optimal transport and the geometry of metric measure spaces.
In his celebrated paper [McC97], McCann studied the convexity of functionals, defined on $\mathcal{P}_{2}(\mathbb{R}^{n})$, with respect to Wasserstein distance. We say that a functional $\mathcal{F}$ is $K$-displacement convex (in the sense of McCann) for some $K\in\mathbb{R}$, if for any $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(\mathbb{R}^{n})$, there is a geodesic ${\mu_{t}}$ in the Wasserstein space, connecting $\mu_{0},\mu_{1}$ such that
$\frac{K}{2}t(1-t)W_{2}^{2}(\mu_{0},\mu_{1})+\mathcal{F}(\mu_{t})\leq(1-t)\mathcal{F}(\mu_{0})+t\mathcal{F}(\mu_{1}).$
Example 1.1 (Potential energy).
Let $f$ be a nonnegative convex function on $\mathbb{R}^{n}$. Then $\mathcal{V}(\mu):=\int_{\mathbb{R}^{n}}f{\mathrm{d}}\mu$ is 00-displacement convex, or displacement convex for simplicity.
There is a family of functionals, called internal energies, which plays crucial role in the study of metric measure spaces. Precisely, the internal energy $\mathcal{H}^{f}_{\mathfrak{m}}$ is defined as
$\mathcal{H}^{f}_{\mathfrak{m}}(\mu):=\left\{\begin{array}[]{ll}\int f(\rho)\,{\mathrm{d}}\mathfrak{m}{}{}{}&\text{if}~{}\mu=\rho\,\mathfrak{m}\\ +\infty&\text{otherwise},\end{array}\right.$
where $f:[0,\infty)\rightarrow[0,\infty]$ is lower semicontinuous, convex and with more than linear growth. When $f=\rho\ln\rho$, $\mathcal{H}^{f}_{\mathfrak{m}}$ is known as Boltzmann entropy, and denoted by ${\rm Ent}_{\mathfrak{m}}$.
Example 1.2 (Boltzmann entropy).
In $(\mathbb{R}^{n},|\cdot|,\mathcal{L}^{n})$, the Boltzmann entropy is displacement convex:
${\rm Ent}_{\mathfrak{m}}(\mu_{t})\leq(1-t){\rm Ent}_{\mathfrak{m}}(\mu_{0})+t{\rm Ent}_{\mathfrak{m}}(\mu_{1}).$
2 Synthetic theory of Ricci curvature
The link between displacement convexity and geometry of metric measure spaces has been deeply studied in the viewpoint of Ricci curvature. On one hand, Cordero-Erausquin, McCann and Schmuckenschläger [CEMS01] show that the Boltzmann entropy is $K$-displacement convex in the Wasserstein space over a compact Riemannian manifold with Ricci curvature bounded from below by $K$. On the other hand, von Renesse and Sturm [vRS05] prove that displacement convexity actually implies Ricci curvature lower bound. Thus, in the setting of Riemannian manifolds, lower Ricci curvature bounds can be (equivalently) characterized by the displacement convexity of certain functionals. This fact is one of the main motivations and starting points of Lott–Sturm–Villani theory about non-smooth metric measure spaces.
Usually, Ricci curvature appears in three places:
  • A:
    estimate of the Jacobian determinant of the exponential map;
  • B:
    Bochner’s formula;
  • C:
    gradient estimate of heat flow.
For example, Cordero-Erausquin–McCann–Schmuckenschläger and von Renesse–Sturm’s results build an connection between these aspects on Riemannian manifolds. These approaches also play crucial roles in the synthetic theory of Ricci curvature (lower bound), and we will recall some related results in this section.
Three curvature-dimension conditions
A: LSV curvature-dimension condition
A curvature-dimension condition for metric measure spaces, called Lott–Sturm–Villani (LSV) curvature-dimension condition today, is introduced independently by Lott–Villani [LV09] and Sturm [Stu06a]. They adopt Cordero-Erausquin– McCann–Schmuckenschläger’s interpolation inequality [CEMS01] as a definition.
Definition 2.1 (${\rm CD}(K,\infty)$ condition).
Let $K\in\mathbb{R}$. A geodesic metric measure space $(X,\mathsf{d},\mathfrak{m})$ verifies ${\rm CD}(K,\infty)$ condition, if, for any two $\mu_{0},\mu_{1}\in\mathcal{P}_{2}(X,\mathsf{d})$, there exists a geodesic $\mu_{t}$ such that for all $t\in(0,1)$:
${\rm Ent}_{\mathfrak{m}}(\mu_{t})\leq(1-t){\rm Ent}_{\mathfrak{m}}(\mu_{0})+t{\rm Ent}_{\mathfrak{m}}(\mu_{1})-\frac{K}{2}(1-t)tW_{2}^{2}(\mu_{0},\mu_{1}).$ (2.1)
B: BE curvature-dimension condition
Barky and Émery [BÉ85] use Bochner’s inequality as a definition of curvature-dimension bounds, in the setting of Dirichlet form theory.
Definition 2.2.
Let $\mathcal{E}$ be a local and symmetric Dirichlet form on the measure space $(X,\mathfrak{m})$ with dense domain ${\rm D}(\mathcal{E})\in L^{2}(X,\mathfrak{m})$ and $(P_{t})_{t\geq 0}$ be the associated Markov semigroup on $L^{2}(X,\mathfrak{m})$ with generator $\Delta$. On a suitable algebra $\mathcal{A}$ of functions dense in the domain ${\rm D}(\mathcal{E})$, one can define the carré du champ operator $\Gamma$ by
$\Gamma(f,g):=\frac{1}{2}\big{(}\Delta(fg)-f\Delta g-g\Delta f\big{)},~{}~{}~{}f,g\in\mathcal{A}.$
Let $\Gamma_{2}$ be a further iteration of the above procedure:
$\Gamma_{2}(f,g):=\frac{1}{2}\big{(}\Delta\Gamma(f,g)-\Gamma(f,\Delta g)-\Gamma(\Delta f,g)\big{)}.$
We say that the measure space $(X,\mathfrak{m})$ verfies ${\rm BE}(K,\infty)$ condition if the following inequality holds:
$\Gamma_{2}(f,f)\geq K\Gamma(f,f).$ (2.2)
Note that, in case $(X,\mathfrak{m})$ is a Riemannian manifold equipped with its volume measure and $\Delta$ is the Laplace–Beltrami operator, the inequality (2.2) is exactly Bochner’s inequality. We refer to [Led14] and [BGL14] for more discussions on this topic.
C: Evolution Variation Inequality (EVI)
In [DS08], Daneri and Savaré provide a different way to study synthetic curvature-dimension condition in non-compact spaces (see Ambrosio–Erbar–Savaré [AES16] for more discussions in the setting of extended metric measure spaces). They show that a stronger formulation of convexity, called EVI-gradient flow (Evolution Variation Inequality), implies (geodesically) convexity. In particular, one can prove McCann’s displacement convexity by showing the existence of EVI-gradient flow in Wasserstein space.
Definition 2.3 (See [AES16]).
Let $(Y,\mathsf{d}_{Y})$ be an extended metric space and $E:Y\rightarrow\mathbb{R}\cup\left\{+\infty\right\}$ be a lower semi-continuous functional, ${\mathrm{D}(E)}:={\left\{E<+\infty\right\}}$. For any $y_{0}\in\mathrm{D}(E)$, we say that $(0,\infty)\ni t\rightarrow y_{t}\in Y$ is an ${\rm EVI}_{K}$ gradient flow of $E$, starting from $y_{0}$, if it is a locally absolutely continuous curve, such that $y_{t}\overset{\mathsf{d}_{Y}}{\rightarrow}y_{0}$ as $t\rightarrow 0$ and the following inequality is satisfied for any $z\in\mathrm{D}(E)$ satisfying $\mathsf{d}_{Y}(z,y_{t})<\infty$ for some (and then all) $t\in(0,\infty)$:
$\frac{1}{2}\frac{{\mathrm{d}}}{{\mathrm{d}}t}\mathsf{d}_{Y}^{2}(y_{t},z)+\frac{K}{2}\mathsf{d}_{Y}^{2}(y_{t},z)\leq E(z)-E(y_{t}),\qquad\text{for a.e. }t\in(0,\infty).$ (2.3)
The $\rm EVI$ gradient flow encodes information not only functional itself but also the structure of the underling space. For example, in a Finsler manifold, the existence of $\rm EVI$ gradient flows of the relative entropy implies that the manifold is Riemannian.
Examples:
The property that the gradient flow of a $K$-convex function $E$ is ${\rm EVI}_{K}$ type, is valid on Euclidean spaces, Hilbert spaces, Alexandrov spaces, ${\rm CAT}$ spaces. For the Wasserstien space $(\mathcal{P}(X),W_{2})$, this property is valid, when $X$ is an Euclidean space, a smooth Riemannian manifold with uniform Ricci lower bound, an Alexandrov space with curvature bounded from below, a Wiener spaces or an $\rm RCD$ space (see §3). We refer the readers to [AGS05, AES16, MS20] for more discussions on this topic. Moreover, we will show its connection with Wasserstein barycenter in Section 3.
In [AGS15], Ambrosio, Gigli and Savaré prove that BE condition, under some natural hypothesis, implies EVI-gradient flow, which then implies a “Riemmanian LSV condition”, the so-called Riemannian curvature-dimension (RCD) condition. Conversely, in [AGS14], they show that $\rm BE$ condition could be deduced from RCD condition. We will discuss this connection in the next subsection.
Riemannian curvature-dimension condition
In [AGS14], Ambrosio, Gigli and Savaré introduce a Riemannian version of the curvature-dimension condition, called Riemannian Curvature-Dimension condition, ruling out Finsler manifolds.
Recall that Cheeger energy (cf. [Che99]) is a functional defined in $L^{2}(X,\mathfrak{m})$ by
${\rm Ch}(f):=\frac{1}{2}\inf\left\{\liminf_{i\rightarrow\infty}\int_{X}\left|{\mathrm{lip}}({f_{i}})\right|^{2}\,{\mathrm{d}}\mathfrak{m}:f_{i}\in{\rm Lip}(X,\mathsf{d}),\left\|f_{i}-f\right\|_{L^{2}}\rightarrow 0\right\},$
where
$\left|{\mathrm{lip}}({h})\right|(x):=\limsup_{y\rightarrow x}\frac{\left|h(y)-h(x)\right|}{\mathsf{d}(x,y)}$
for $h\in\mathop{\rm Lip}\nolimits(X,\mathsf{d})$. We define the Sobolev space $W^{1,2}(X,\mathsf{d},\mathfrak{m})$ by
$W^{1,2}(X,\mathsf{d},\mathfrak{m}):=\left\{f\in L^{2}(X,\mathfrak{m}):{\rm Ch}(f)<\infty\right\}.$
Definition 2.4 (${\rm RCD}(K,\infty)$ condition).
Let $K\in\mathbb{R}$. A metric measure space $(X,\mathsf{d},\mathfrak{m})$ verifies Riemannian curvature bounded from below by $K$ (or $(X,\mathsf{d},\mathfrak{m})$ is an ${\rm RCD}(K,\infty)$ space) if it satisfies ${\rm CD}(K,\infty)$ condition and the Cheeger energy ${\rm Ch}$ is a quadratic form (Riemannian structure) in the sense that
${\rm Ch}(f+g)+{\rm Ch}(f-g)=2{\rm Ch}(f)+2{\rm Ch}(g)\qquad\text{for all}~{}f,g\in W^{1,2}(X).$ (2.4)
Formally speaking, an $\rm RCD$ space is a $\rm CD$ space equipped with a Riemannian structure. Remarkably, the above $\rm RCD$ condition can be equivalently defined via gradient flow.
Theorem 2.5 ([AGS14], Theorem 5.1; [AGMR15], Theorem 6.1).
Let $(X,\mathsf{d},\mathfrak{m})$ be a metric measure space and $K\in\mathbb{R}$. Then $(X,\mathsf{d},\mathfrak{m})$ satisfies ${\rm RCD}(K,\infty)$ condition if only and if $(X,\mathsf{d},\mathfrak{m})$ is a length space satisfying an exponential growth condition
$\int_{X}e^{-c\mathsf{d}(x_{0},x)^{2}}\,{\mathrm{d}}\mathfrak{m}(x)<\infty,\qquad\text{for all }x_{0}\in X\text{ and }c>0,$
and for all $\mu\in\mathcal{P}_{2}(X,\mathsf{d})$ there exists an ${\rm EVI}_{K}$ gradient flow, in $(\mathcal{P}_{2}(X,\mathsf{d}),W_{2})$, of ${\rm Ent}_{\mathfrak{m}}$ starting from $\mu$.
3 Wasserstein barycenter
In this section we review some recent results obtained by the authors in [HLZ24] concerning Wasserstein barycenter problem.
Let $\Omega$ be a Borel probability measure on an extended metric space $(Y,\mathsf{d}_{Y})$, a barycenter of $\Omega$ is defined as a minimizer of
$x\mapsto\int_{Y}\mathsf{d}_{Y}^{2}(x,y)\,{\mathrm{d}}\Omega(y)$
and the variance of $\Omega$ is defined as
${\rm Var}(\Omega):=\inf_{x\in X}\int_{Y}\mathsf{d}_{Y}^{2}(x,y)\,{\mathrm{d}}\Omega(y).$
When $(Y,\mathsf{d}_{Y})$ is a Wasserstein space, a barycenter of $\Omega$ is also called Wasserstein barycenter. This kind of problems, called Wasserstein barycenter problems, draw particular interests in optimal transport and its related fields, as it gives a natural but non-linear way to interpolate between a distribution of measures.
In metric measure geometry setting, Agueh–Carlier [AC11] in Euclidean spaces, Kim–Pass [KP17] and Ma [Ma23] in Riemannian manifolds, Jiang [Jia17] in Alexandrov spaces, establish the existence and uniqueness of Wasserstein barycenter, and the absolutely continuity of the Wasserstein barycenter with respect to the canonical reference measures in these cases. As a consequence, they get the following Wasserstein Jensen’s inequality, which characterizes the convexity of certain functionals via Wasserstein barycenter:
For any $K$-displacement convex functional $\mathcal{F}:\mathcal{P}_{2}(X,\mathsf{d})\to\mathbb{R}\cup\{+\infty\}$ in the sense of McCann [McC97], a probability measure $\Omega$ on $\mathcal{P}_{2}(X,\mathsf{d})$ and its unique barycenter $\bar{\mu}$, it holds
$\mathcal{F}(\bar{\mu})\leq\int_{\mathcal{P}_{2}(X,\mathsf{d})}\mathcal{F}(\mu)\,{\mathrm{d}}\Omega(\mu)-\frac{K}{2}\int_{\mathcal{P}_{2}(X,\mathsf{d})}W_{2}^{2}(\bar{\mu},\mu)\,{\mathrm{d}}\Omega(\mu).$ (WJI)
The above inequality can be understood as a general version of “displacement convexity” with respect to Wasserstein barycenter. Motivated by the synthetic theory of Ricci curvature, it is natural to ask:
Does the Wasserstein Jensen’s inequality hold on $\rm RCD$ and $\rm CD$ spaces?
One may try to imitate the approach used in $\mathbb{R}^{n}$, manifolds and Alexandrov spaces. A barrier lies in the lack of sectional curvature bound, which plays an important role in deriving the regularity of Wasserstein barycenter in their proofs. To overcome this difficulty, we deal with Wasserstein barycenter in a synthetic approach, by making full use of Jensen’s inequality. We observe that the Wasserstein Jensen’s inequality actually encodes the regularity of barycenter. For example, applying $\mathcal{F}={\rm Ent}_{\mathfrak{m}}$ in (3.1), we get
${\rm Ent}_{\mathfrak{m}}(\bar{\mu})\leq\int_{\mathcal{P}_{2}(X,\mathsf{d})}{\rm Ent}_{\mathfrak{m}}(\mu)\,{\mathrm{d}}\Omega(\mu)-\frac{K}{2}\int_{\mathcal{P}_{2}(X,\mathsf{d})}W_{2}^{2}(\bar{\mu},\mu)\,{\mathrm{d}}\Omega(\mu).$ (3.1)
So ${\rm Ent}_{\mathfrak{m}}(\bar{\mu})<+\infty$ in case $\int_{\mathcal{P}_{2}(X,\mathsf{d})}{\rm Ent}_{\mathfrak{m}}(\mu)\,{\mathrm{d}}\Omega(\mu)<+\infty$ and $\Omega$ has finite variance.
Motivated by the work of Daneri and Savaré [DS08], we realize that the existence of $\rm EVI$ gradient flow, introduced in Section 2, implies the Jensen’s inequality.
Theorem 3.1 ($\rm EVI$ implies Jensen’s inequality).
Let $(Y,\mathsf{d}_{Y})$ be an extended metric space, $K\in\mathbb{R}$, $E$ be as in Definition 2.3 and $\mu$ be a probability measure over $Y$ with finite variance. Let $\epsilon\geq 0.$ Assume that ${\rm EVI}_{K}$ gradient flow of $E$ exists for some initial data $y\in Y$ satisfying
$\int_{Y}\mathsf{d}_{Y}^{2}(y,z)\,{\mathrm{d}}\mu(z)\leq{\rm Var}(\mu)+\epsilon.$ (3.2)
Then the following inequality holds:
$E(y_{t})\leq\int_{Y}E(z)\,{\mathrm{d}}\mu(z)-\frac{K}{2}{\rm Var}(\mu)+\frac{\epsilon}{2I_{K}(t)},$ (3.3)
where $y_{t}$ is the ${\rm EVI}_{K}$ gradient flow of E starting from $y$. In particular, if $\mu$ has a barycenter $\bar{y}$ and there is an ${\rm EVI}_{K}$ gradient flow of $E$ starting from $\bar{y}$, then we have Jensen’s inequality:
$E(\bar{y})\leq\int_{Y}E(z)\,{\mathrm{d}}\mu(z)-\frac{K}{2}\int_{Y}\mathsf{d}^{2}(\bar{y},z)\,{\mathrm{d}}\mu(z).$ (3.4)
The following proposition, proved in [DS08, Proposition 3.1] (see also [AES16]), provides an integral version of ${\rm EVI}$ type gradient flows. We will use this formulation to prove Jensen’s inequality.
Proposition 3.2 (Integral version of $\rm EVI$).
Let $E$, $K$ and $(y_{t})_{t>0}$ be as in Definition 2.3, then $(y_{t})_{t>0}$ is an ${\rm EVI}_{K}$ gradient flow if and only if it satisfies
$\frac{e^{K(t-s)}}{2}\mathsf{d}_{Y}^{2}(y_{t},z)-\frac{1}{2}\mathsf{d}_{Y}^{2}(y_{s},z)\leq I_{K}(t-s)\big{(}E(z)-E(y_{t})\big{)},\quad\forall 0\leq s\leq t,$ (3.5)
where $I_{K}(t)$:=$\int_{0}^{t}e^{Kr}\,{\mathrm{d}}r$.
Proof of Theorem 3.1.
Integrating (3.5) in $z$ with respect to $\mu$ and choosing $s=0$, we get
$\frac{e^{Kt}}{2}\int_{Y}\mathsf{d}_{Y}^{2}(y_{t},z)\,{\mathrm{d}}\mu(z)-\frac{1}{2}\int_{Y}\mathsf{d}_{Y}^{2}(y,z)\,{\mathrm{d}}\mu(z)\leq I_{K}(t)\left(\int_{Y}E(z)\,{\mathrm{d}}\mu(z)-E(y_{t})\right).$ (3.6)
From the definition of barycenter, we can see that $\int_{Y}\mathsf{d}_{Y}^{2}(y_{t},z)\,{\mathrm{d}}\mu(z)\geq{\rm Var}(\mu)$. Combining with (3.2) and (3.6), we get
$\frac{e^{Kt}-1}{2}{\rm Var}(\mu)-\frac{\epsilon}{2}\leq I_{K}(t)\left(\int_{Y}E(z)\,{\mathrm{d}}\mu(z)-E(y_{t})\right).$ (3.7)
Dividing both sides of (3.7) by $I_{K}(t)$, we get (3.3). In particular, if $y$ is a barycenter of $\mu$, we can take $\epsilon=0$ in (3.3), so that
$E(y_{t})\leq\int_{Y}E(z)\,{\mathrm{d}}\mu(z)-\frac{K}{2}{\rm Var}(\mu).$
Letting $t\rightarrow 0$, and recalling the lower semi-continuity of $E$ and $y_{t}\overset{\mathsf{d}_{Y}}{\rightarrow}y_{0}$, we get Jensen’s inequality (3.4). ∎
Combining Theorem 2.5 and Theorem 3.1 we establish the Wasserstein Jensen’s inequality in $\rm RCD$ spaces.
Example 3.3.
Let $(X,\mathsf{d},\mathfrak{m})$ be an ${\rm RCD}(K,\infty)$ space. Let $\Omega$ be a probability measure over $\mathcal{P}_{2}(X,\mathsf{d})$ with finite variance. Then for any barycenter $\bar{\mu}$ of $\Omega$, it holds
${\rm Ent}_{\mathfrak{m}}(\bar{\mu})\leq\int_{\mathcal{P}_{2}(X,\mathsf{d})}{\rm Ent}_{\mathfrak{m}}(\mu)\,{\mathrm{d}}\Omega(\mu)-\frac{K}{2}\int_{\mathcal{P}_{2}(X,\mathsf{d})}W_{2}^{2}(\bar{\mu},\mu)\,{\mathrm{d}}\Omega(\mu).$ (3.8)
Furthermore, the geometry of some extended measure spaces, such as Wiener spaces [AES16] and configuration spaces [EH15], can also be characterized via $\rm EVI$ gradient flows with initial data in ${\rm D}({\rm Ent})$. Concerning the application of Theorem 3.1 for extended metric measure spaces, it is worth mentioning that one should take care of the existence of $\rm EVI$ gradient flows. In general, this is a highly non-trivial problem.
Example 3.4.
Let $(X,H,\gamma)$ be an abstract Wiener space equipped with the Cameron–Martin distance $\mathsf{d}_{H}$ and the Gaussian measure $\gamma$. Let $\Omega$ be a probability measure over $\mathcal{P}(X)$ with finite variance.
Then for any barycenter $\bar{\mu}$ of $\Omega$ it holds
${\rm Ent}_{\gamma}(\bar{\mu})\leq\int_{\mathcal{P}(X)}{\rm Ent}_{\gamma}(\mu)\,{\mathrm{d}}\Omega(\mu)-\frac{1}{2}\int_{\mathcal{P}(X)}W_{2}^{2}(\bar{\mu},\mu)\,{\mathrm{d}}\Omega(\mu).$ (3.9)
Example 3.5 (See [EH15]).
Let $(\mathbf{Y},\mathsf{d}_{\mathbf{Y}},\pi)$ be an extended metric measure space consisting of the configuration space $\mathbf{Y}$ over a Riemannian manifold $M$, the intrinsic metric $\mathsf{d}_{\mathbf{Y}}$ and the Poisson measure $\pi$. Assume that $M$ has Ricci curvature bounded below by $K\in\mathbb{R}$.
Then for any probability measure $\Omega$ over $\mathcal{P}(\mathbf{Y})$ with finite variance, its barycenter $\bar{\mu}$ satisfies
${\rm Ent}_{\pi}(\bar{\mu})\leq\int_{\mathcal{P}(X)}{\rm Ent}_{\pi}(\mu)\,{\mathrm{d}}\Omega(\mu)-\frac{K}{2}\int_{\mathcal{P}(X)}W_{2}^{2}(\bar{\mu},\mu)\,{\mathrm{d}}\Omega(\mu).$
Remark 3.6.
Concerning general $\rm CD$ spaces, it seems hard to prove Wasserstein Jensen’s inequality. At the moment, we even do not know whether this is true for Finsler manifolds or not.
4 Barycenter Curvature-Dimension condition
We have seen the links between displacement convexity, Wasserstein barycenter (Jensen’s inequality) and curvature-dimension conditions of metric measure spaces. This motivates us to propose a new curvature-dimension condition, which covers some non-trivial extended metric measure spaces (also non-geodesic), e.g. Wiener spaces and Configuration spaces over a Riemannian manifold.
Since barycenter plays a crucial role, we call this new curvature-dimension condition, Barycenter Curvature-Dimension condition and denote it by BCD.
Definition 4.1 (${\rm BCD}(K,\infty)$ condition).
Let $K\in\mathbb{R}$. We say that an extended metric measure space $(X,\mathsf{d},\mathfrak{m})$ verifies ${\rm BCD}(K,\infty)$ condition, if for any probability measure $\Omega\in\mathcal{P}_{2}(\mathcal{P}(X),W_{2})$, concentrated on finitely many measures, there exists a barycenter $\bar{\mu}$ of $\Omega$ such that the following Jensen’s inequality holds:
${\rm Ent}_{\mathfrak{m}}(\bar{\mu})\leq\int_{\mathcal{P}(X)}{\rm Ent}_{\mathfrak{m}}(\mu)\,{\mathrm{d}}\Omega(\mu)-\frac{K}{2}{\rm Var}(\Omega).$ (4.1)
Remark 4.2.
If we take $\Omega=(1-t)\delta_{\mu_{0}}+t\delta_{\mu_{1}}$ and $(X,\mathsf{d})$ is a geodesic space, (4.1) implies the Lott–Sturm–Villani curvature-dimension condition.
Remark 4.3.
Similar to Lott–Sturm–Villani’s ${\rm CD}(K,N)$ condition (cf. [LV09, Stu06b]) and Bakry–Émery’s ${\rm BE}(K,N)$ (cf. [BGL14]), there is a finite dimensional version of $\rm BCD$ condition, ${\rm BCD}(K,N)$ condition; see [HLZ24] for details.
Not surprisingly, we can establish that the moduli space of compact metric measure spaces satisfying barycenter curvature-dimension condition, is closed under measured Gromov–Hausdorff convergence.
Theorem 4.4 (Stability in measured Gromov–Hausdorff topology).
Let $\left\{(X_{i},\mathsf{d}_{i},\nu_{i})\right\}_{i=1}^{\infty}$ be a sequence of compact ${\rm BCD}(K,\infty)$ metric measure spaces with $K\in\mathbb{R}$. If $(X_{i},\mathsf{d}_{i},\nu_{i})$ converges to $(X,\mathsf{d},\nu)$ in the measured Gromov–Hausdorff sense as $n\rightarrow\infty$, then $(X,\mathsf{d},\nu)$ is also a ${\rm BCD}(K,\infty)$ space.
As explained in Section 2, the difference between $\rm CD$ condition and $\rm RCD$ condition lies in whether they have a “Riemannian structure” or not. It is natural to ask:
Problem 1:
Do Finsler manifolds with lower Ricci curvature bound from below satisfy BCD condition? Is there a “Riemmanian Barycenter-curvature dimension condition”?
Problem 2:
What is the difference between $\rm CD$ condition and $\rm BCD$ condition?
Since spaces satisfying $\rm BCD$ condition, we have discovered, are geodesic or “almost geodesic”, it is interesting to find a $\rm BCD$ space which is not geodesic.
Problem 3:
Is there a non-geodesic $\rm BCD$ space?
At the end of this paper, we introduce some new geometric inequalities, as simple but interesting applications of our BCD theory. For simplicity, we will only deal with ${\rm BCD}$ metric measure spaces. We also refer the readers to a recent paper [KW22] by Kolesnikov and Werner for more applications of Wasserstein barycenter.
Proposition 4.5 (Multi-marginal logarithmic Brunn–Minkowski inequality).
Let $(X,\mathsf{d},\mathfrak{m})$ be a ${\rm BCD}(0,\infty)$ metric measure space, $E_{1},...,E_{n}$ be bounded measurable sets with positive measure and $\lambda_{1},\dots,\lambda_{n}\in(0,1)$ with $\sum_{i}\lambda_{i}=1$. Then
$\mathfrak{m}(E)\geq\mathfrak{m}(E_{1})^{\lambda_{1}}...\mathfrak{m}(E_{n})^{\lambda_{n}}$
where
$E:=\left\{x~{}\text{is the barycenter of}~{}\sum_{i=1}^{n}\lambda_{i}\delta_{x_{i}}:x_{i}\in E_{i},i=1,\dots,n\right\}.$
Proposition 4.6 (A functional Blaschke–Santaló type inequality).
Let $(X,\mathsf{d},\mathfrak{m})$ be a ${\rm BCD}(1,\infty)$ metric measure space. Then we have
$\prod_{i=1}^{k}\int_{X}e^{f_{i}}\,{\mathrm{d}}\mathfrak{m}\leq 1$
for any measurable functions $f_{i}$ on $X$ such that $\frac{e^{f_{i}}}{\int e^{f_{i}}\,{\mathrm{d}}\mathfrak{m}}\in\mathcal{P}_{2}(X,\mathsf{d})$ and
$\sum_{i=1}^{k}f_{i}(x_{i})\leq\frac{1}{2}\inf_{x\in X}\sum_{i=1}^{k}{\mathrm{d}}(x,x_{i})^{2}\qquad\forall x_{i}\in X,i=1,2,...,k.$
Declaration. The authors declare that there is no conflict of interest and the manuscript has no associated data.
Acknowledgement: This work is supported in part by Young Scientist Programs of the Ministry of Science and Technology of China (2021YFA1000900, 2021YFA1002200), and National Natural Science Foundation of China (12201596).